karlieeeeeee:

Is (–6, –9) a solution to this system of equations? 13x − 8y = –6 y = –9

simsharrison:

Take one of the equations and solve for one of the variables. For example, take 3x-4y=2 --> 3x = 4y+2 --> x = (4y+2)/3

simsharrison:

@jhonyy9 ? ^ is that on the right track?

simsharrison:

@karlieeeeeee

jhonyy9:

just substitute these value of x and y inside this equation make th calcule and look what will get

simsharrison:

Oh so we will multiply -6 x 13 =?

simsharrison:

all i did was replace -6 with x

jhonyy9:

13x -8y = -6 13*(-6) -8*(-9) = -6 so this is right ?

jhonyy9:

-78 +72 = ?

jhonyy9:

=-6 so -6 = -6 what result from these ?

simsharrison:

1

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 no result these values of x and y are solutions of this system of equations \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

ok ?

simsharrison:

yes

jhonyy9:

good